# HERBench: A Benchmark for Multi-Evidence Integration in Video Question Answering

### v1

@@ 4 Experiments
Model TR&C R&T GC&V ME&N Overall Avg. TSO MPDR ASII Avg. AGBI AGAR AGLT Avg. FAM SVA FOM Avg. MEGL AC RLPC Avg. GPT-4.1 [openai_gpt41_2025] 18.9 29.7 27.7 25.4 78.0 59.1 61.0 66.0 30.4 38.9 41.9 37.1 25.5 24.3 37.3 29.0 39.4 Gemini-2.5-Flash [gemini25_2025] 28.6 35.8 24.8 29.7 75.2 71.4 63.1 69.9 29.2 31.3 44.2 34.9 22.6 26.6 31.2 26.8 40.3 Qwen2.5-VL-72B [qwen25vl_72b] 10.4 42.6 27.8 26.9 74.4 76.1 62.2 70.9 25.6 50.6 33.5 36.6 18.1 23.0 32.0 24.4 39.7 Gemma-3-27B [gemma3_27b_2025] 38.4 42.0 15.7 32.0 69.0 50.5 55.6 58.4 21.8 14.3 28.4 21.5 15.7 29.0 25.7 23.5 33.8 LLaMA-4-Scout-17B [llama4_scout_17b_2025] 6.2 30.0 20.1 18.8 64.7 51.6 55.6 57.3 19.3 36.5 20.7 25.5 17.2 26.1 29.4 24.2 31.4 InternVL3.5-14B [internvl35_8b_2025] 43.9 38.8 30.3 37.7 75.9 69.4 62.6 69.3 26.8 22.8 43.8 31.1 25.3 20.8 37.3 27.8 41.5 Ovis-2.5-9B [ovis25_9b_2025] 0.1 30.6 26.0 18.9 79.7 76.2 64.7 73.5 33.6 57.2 49.6 46.8 27.5 23.4 36.7 29.2 42.1 InternVL3.5-8B [internvl35_8b_2025] 41.3 31.3 28.1 33.6 77.6 71.6 61.4 70.2 26.3 21.2 41.5 29.7 33.1 21.2 38.1 30.8 41.1 LLaVA-OneVision1.5-8B [llava_onevision15_8b_2025] 26.6 28.7 23.0 26.1 76.8 67.5 58.8 67.7 29.8 33.9 37.1 33.6 25.2 17.6 31.9 24.9 38.1 Qwen3-VL-8B [qwen3vl_8b_2025] 2.2 28.7 26.0 19.0 74.6 69.6 61.9 68.7 30.0 51.3 40.4 40.6 18.8 21.8 34.9 25.2 38.3 MiniCPM-V4.5-8B [minicpm_v45_8b_2025] 19.1 26.3 26.0 23.8 77.9 72.3 63.2 71.1 30.2 43.7 45.2 39.7 24.1 22.9 27.9 24.9 39.9 Qwen2.5-VL-7B [qwen25vl_72b] 14.6 28.0 22.9 21.8 69.7 59.3 52.9 60.6 33.0 36.0 47.1 38.7 21.1 20.3 26.3 22.6 35.9 LLaVA-OneVision-7B [llava_onevision_7b_2024] 33.3 24.9 23.7 27.3 67.1 58.0 52.3 59.1 28.9 22.4 38.9 30.1 22.8 22.4 32.8 26.0 35.6 Avg. 23.1 31.9 25.5 26.8 74.5 66.3 59.7 66.8 28.4 35.2 40.9 34.8 23.2 23.0 32.7 26.3 38.2
Table 2: Main results on the HERBench. We report per-task accuracy (%) for 13 leading MLLMs. The highest performance in each task column is marked in bold. The 4 largest-size models (first rows) were run on a representative 10% subset ($\sim$2.6K questions), while the remaining models were evaluated on the full benchmark.
@@
We evaluate 13 prominent MLLMs, spanning closed-source models (GPT-4.1 [openai_gpt41_2025], Gemini-2.5-Flash [gemini25_2025]) and a diverse range of open-source systems, detailed in Table 2. This selection allows for a broad assessment of state-of-the-art capabilities across scales and design paradigms. To isolate reasoning ability from evidence retrieval, we standardize the visual input. All models receive an identical budget of 16 frames, sampled uniformly across the video. This fixed input ensures that performance differences are attributable to the models’ multi-evidence integration capabilities, not to varied frame selection. We report top-1 accuracy.
@@ Results.
As shown in Table 2, performance is systematically poor. The mean accuracy across all 13 state-of-the-art models is 38.2%, with the best model (Ovis-2.5-9B [ovis25_9b_2025]) reaching only 42.1% and the lowest (LLaMA-4-Scout-17B [llama4_scout_17b_2025]) at 31.4%. This narrow performance band, just 11–22 percentage above the 20% random baseline, reveals that failure to integrate dispersed evidence is a pervasive limitation across all current architectures.
@@ 5.1 Isolating the Evidence Retrieval Bottleneck
To address (Q1), i.e. the impact of evidence retrieval, we compare five strategies (all operating in the same BLIP [li2022blip] embedding space for fairness): AKS [tang2025aks] learns a keyframe policy that balances relevance and temporal coverage; BOLT-ITS [liu2025bolt] using inverse transform sampling to select query-relevant frames; Uniform takes evenly spaced frames; Vanilla-BLIP retrieves frames with highest cosine similarity to the question; Oracle Frames (OF) use frame indices gathered from our benchmark’s construction pipeline along with a few complementary non-evidence frames. They are applied only in tasks where relevant evidence is scarce or confined to very short portions of the video (notably [TSO], [FAM], [SVA]).  [^1] We use “OF” to denote the curated set of evidence frame indices plus a few additional non-evidence frames. The purpose of the additional frames is to fill the fixed frame budget (e.g., 16 frames), making OF a ”best-case” selection strategy that is comparable to AKS and BOLT-ITS. This setup tests if a model can identify the correct evidence when it is guaranteed to be retrieved, and is distinct from the oracle-only analysis in Sec. 5.2, which tests pure fusion. Curation exists only for a subset of tasks, so its effect is not uniform across the benchmark.
@@ 5.2 Evidence Aggregation with Oracle-Only Frames
Having established that evidence retrieval is a significant bottleneck (Sec. 5.1), we now turn to (Q2): can models effectively aggregate evidence even when retrieval uncertainty is removed? To isolate the fusion capability from the retrieval challenge, we conduct a targeted study on a subset of HERBench supplying models with only the manually curated ground-truth frames (the ”oracle” set).
Figure 5: Top-1 frame share under oracle-only frames. Violin/box plots show the distribution of the maximum normalized frame-importance share across oracle-only frames for three models (InternVL3.5-14B, Ovis-2.5, Qwen3-VL-8B), split by Correct vs. Incorrect predictions. For each item, we compute leave-one-out deltas of the log-probability of the model’s predicted option and normalize them to per-frame shares; the plotted statistic is the largest share (Top-1). Correct predictions allocate credit more evenly across frames (typically $\sim$0.5), whereas errors over-concentrate on a single frame (often $\sim 0.8$), indicating insufficient multi-evidence fusion even when only evidence-bearing frames are provided.
@@ 6 Conclusion
We introduce HERBench, a VideoQA benchmark comprising 26,806 questions across 12 tasks, each constructed to structurally enforce aggregating $k\geq 3$ distinct, temporally separated visual cues. To quantify evidential demand, we propose the Minimum Required Frame-Set (MRFS) metric. Cross-benchmark validation confirms HERBench imposes substantially higher evidential requirements (mean MRFS 5.5 vs. 2.6-4.2 on prior datasets). Evaluating 13 state-of-the-art MLLMs reveals a limited accuracy range (31.4–42.1%), only slightly above the 20% chance-level baseline (corresponding to randomly selecting one of five answer options), exposing pervasive limitations. We identify two core bottlenecks: (i) a retrieval deficit, as frame selectors fail to access all necessary cues, and (ii) a fusion deficit, where models fail to integrate evidence even when provided. This fusion failure manifests as over-concentration on single frames. Our analysis thus isolates specific, actionable deficits in aggregation and selection, providing a clear target for improving future Video-LLMs. This benchmark will thus serve as a critical tool to guide the development of next-generation models capable of genuine compositional reasoning. By making multi-evidence aggregation both unavoidable and quantifiable, HERBench establishes a principled target for advancing compositional video understanding and exposes clear headroom for progress beyond single-cue success.
HERBench: A Benchmark for Multi-Evidence Integration in Video Question Answering
@@ 7.1 Track Ranking and Selection
We utilize a detection-tracking stack where RF-DETR detectors feed into the DeepSORT multi-object tracker. We apply a high-recall detector with a confidence threshold of 0.3 and a per-frame cap of 300 detections. Association uses a two-stage IoU matching: high-confidence detections (score >0.5>0.5) are matched with an IoU threshold of 0.7, followed by lower-confidence detections with a relaxed IoU threshold of 0.35.
@@
Prior to ranking, we enforce a hard filter: we keep only the COCO “person” class, require length $\geq 20$ frames, average area $\geq 5,500$ pixels, and require the track center to fall within the central safe region (frame cropped by 10% margins) in at least 5 frames.
Figure 6: Tracking, post-processing, and ranking pipeline. RF-DETR detections are linked with DeepSORT into raw person tracks, followed by outlier removal, gap interpolation, and Gaussian smoothing. A TrackRanker then scores and selects salient trajectories, which are passed to an MLLM descriptor module to generate temporally decoupled appearance (A) and behavior (B) cards that serve as the scaffold for downstream HERBench tasks.
@@ 7.2 Decoupled Descriptor Generation
For each selected track, we generate disentangled descriptions using GPT-4o. We sample 10-11 crops, reserving the first and last 20% of the trajectory for Appearance (A-cards) and the middle 60% for Behavior (B-cards). This ensures a temporal gap of at least 30 frames between appearance and behavior cues. An example of the resulting disentangled A- and B-cards for a single track is shown in Figure 7. We use the following prompt structure:
@@ 7.3 Spatial Operations and Region Definitions
For [RLPC], we also define rectangular ROIs (e.g., frame halves or specific zones). A track is counted as visiting an ROI if, at any frame, at least 50% of its bounding box area lies within the region (Intersection-Over-Box $\geq 0.5$). We count the unique track IDs that satisfy this predicate to derive people counts under spatial constraints.
Duration computation (MPDR).
@@ 7.4 Scene Card Perturbations
We use TransNetV2 for shot boundary detection. For the Scene Verification Arrangement (SVA) task, faithful scene cards are generated via an MLLM using the following prompt:
@@
To generate negative samples for SVA, we prompt the model to modify faithful descriptions by altering 2-5 atomic details. The prompt constraints ensure:
@@
An example of a faithful scene card and its perturbed counterpart used for the SVA task is shown in Figure 8.
@@ 7.5 Corpus-Plausible Foil Generation
  • False Action Memory (FAM): We sample a “false” action by pairing an object present in the video with an action from the corpus that does not occur in the current video.
  • False Object Memory (FOM): We select an absent object from the corpus-wide index that is compatible with actions present in the video (e.g., if “cutting” occurs, “carrot” is a valid distractor if absent).
  • Action Counting (AC): Distractor counts are generated such that the correct count’s rank varies uniformly across options.
  • Action Sequence Integrity (ASII): We sample a 5-event ground-truth timeline. Distractors are generated using two perturbation functions: swap_mid (swapping two non-adjacent events) and rotate (shifting the sequence). Crucially, we verify against the event log that the perturbed timeline does not accidentally exist in the video.
@@ 7.6 Text-Only Bias Filtering Details
To suppress language priors, we apply a rigorous Text-Only Filtering stage. We discard any question correctly answered by $\geq 3$ of 4 blind LLMs (Qwen2-7B, Qwen2.5-7B, Llama-3-8B, and Vicuna-7B v1.5). This step rejects approximately 10% of candidates (e.g., questions answerable via object-color co-occurrence priors).
@@ 7.7 Human Verification Protocol
Verification checklist.
Experts conduct verification on a stratified 15% sample. The checklist includes:
  • Minimum Frame-Set: Confirming the question requires $k\geq 3$ distinct frames.
  • Uniqueness: Ensuring a unique, objective ground-truth answer exists.
  • Disentanglement: Verifying A/B cards do not leak information.
This process resulted in an 17.8% rejection rate.

### v2

@@ 4 Experiments
Model TR&C R&T GC&V ME&N Overall Avg. TSO MPDR ASII Avg. AGBI AGAR AGLT Avg. FAM SVA FOM Avg. MEGL AC RLPC Avg. GPT-4.1 [34] 18.9 29.7 27.7 25.4 78.0 59.1 61.0 66.0 30.4 38.9 41.9 37.1 25.5 24.3 37.3 29.0 39.4 Gemini-2.5-Flash [7] 28.6 35.8 24.8 29.7 75.2 71.4 63.1 69.9 29.2 31.3 44.2 34.9 22.6 26.6 31.2 26.8 40.3 Qwen2.5-VL-72B [3] 10.4 42.6 27.8 26.9 74.4 76.1 62.2 70.9 25.6 50.6 33.5 36.6 18.1 23.0 32.0 24.4 39.7 Gemma-3-27B [13] 38.4 42.0 15.7 32.0 69.0 50.5 55.6 58.4 21.8 14.3 28.4 21.5 15.7 29.0 25.7 23.5 33.8 LLaMA-4-Scout-17B [31] 6.2 30.0 20.1 18.8 64.7 51.6 55.6 57.3 19.3 36.5 20.7 25.5 17.2 26.1 29.4 24.2 31.4 InternVL3.5-14B [45] 43.9 38.8 30.3 37.7 75.9 69.4 62.6 69.3 26.8 22.8 43.8 31.1 25.3 20.8 37.3 27.8 41.5 Ovis-2.5-9B [27] 0.1 30.6 26.0 18.9 79.7 76.2 64.7 73.5 33.6 57.2 49.6 46.8 27.5 23.4 36.7 29.2 42.1 InternVL3.5-8B [45] 41.3 31.3 28.1 33.6 77.6 71.6 61.4 70.2 26.3 21.2 41.5 29.7 33.1 21.2 38.1 30.8 41.1 LLaVA-OneVision1.5-8B [1] 26.6 28.7 23.0 26.1 76.8 67.5 58.8 67.7 29.8 33.9 37.1 33.6 25.2 17.6 31.9 24.9 38.1 Qwen3-VL-8B [2] 2.2 28.7 26.0 19.0 74.6 69.6 61.9 68.7 30.0 51.3 40.4 40.6 18.8 21.8 34.9 25.2 38.3 MiniCPM-V4.5-8B [56] 19.1 26.3 26.0 23.8 77.9 72.3 63.2 71.1 30.2 43.7 45.2 39.7 24.1 22.9 27.9 24.9 39.9 Qwen2.5-VL-7B [3] 14.6 28.0 22.9 21.8 69.7 59.3 52.9 60.6 33.0 36.0 47.1 38.7 21.1 20.3 26.3 22.6 35.9 LLaVA-OneVision-7B [19] 33.3 24.9 23.7 27.3 67.1 58.0 52.3 59.1 28.9 22.4 38.9 30.1 22.8 22.4 32.8 26.0 35.6 Avg. 23.1 31.9 25.5 26.8 74.5 66.3 59.7 66.8 28.4 35.2 40.9 34.8 23.2 23.0 32.7 26.3 38.2
Table 2: HERBench results. We report per-task accuracy (%) for 13 leading MLLMs. The highest performance in each task column is marked in bold. The 4 largest-size models (first rows) were run on a representative 10% subset ($\sim$2.6K questions), while the remaining models were evaluated on the full benchmark.
@@
We evaluate 13 prominent MLLMs, including both closed- and open-source systems (Table 2). To isolate the effects of evidence aggregation from frame selection, all models receive the same budget of 16 uniformly sampled frames from each video.
@@ Results.
As shown in Table 2, performance is systematically poor. The mean accuracy across all 13 state-of-the-art models is 38.2%, with the best model (Ovis-2.5-9B [27]) reaching only 42.1% and the lowest (LLaMA-4-Scout-17B [31]) at 31.4%. This narrow performance band, just 11–22 percentage above the 20% random baseline, reveals that failure to integrate dispersed evidence is a pervasive limitation across all current architectures.
@@ 5.1 Isolating the Evidence Retrieval Bottleneck
To address (Q1), i.e. the impact of evidence retrieval, we compare five strategies (all operating in the same BLIP [21] embedding space for fairness): AKS [43] learns a keyframe policy that balances relevance and temporal coverage; BOLT-ITS [25] using inverse transform sampling to select query-relevant frames; Uniform takes evenly spaced frames; Vanilla-BLIP retrieves frames with highest cosine similarity to the question; Oracle Frames (OF) use frame indices gathered from our benchmark’s construction pipeline along with a few non-evidence frames to match the fixed frame budget (they are meant to represent a best-case retrieval setting rather than the oracle-only fusion setting studied in Sec. 5.2). They are applied only in tasks where relevant evidence is scarce or confined to very short portions of the video (notably [TSO], [FAM], [SVA]).
@@ 5.2 Evidence Aggregation with Oracle-Only Frames
Having established that evidence retrieval is a significant bottleneck (Sec. 5.1), we now turn to (Q2): can models effectively aggregate evidence even when retrieval uncertainty is removed? To isolate the fusion capability from the retrieval challenge, we conduct a targeted study on a subset of HERBench supplying models with only the manually curated ground-truth frames (the ”oracle” set). In a parallel human study under the same oracle presentation format, annotators achieved 95.7% accuracy, indicating that the curated oracle frame-set is generally sufficient for resolving the question.
Figure 5: Top-1 frame share under oracle-only frames. We plot the maximum normalized frame-importance share for three models, separated by correct and incorrect predictions. Correct answers distribute importance more evenly, whereas errors concentrate it on a single frame, indicating weak multi-evidence fusion even when evidence-bearing frames are provided.
@@ 6 Conclusion
We introduced HERBench, a VideoQA benchmark comprising 26,806 questions across 12 tasks, each designed to structurally enforce the aggregation of $k\geq 3$ distinct, temporally separated visual cues. To quantify evidential demand, we proposed the Minimum Required Frame-Set (MRFS) metric; under a canonical evaluation protocol, HERBench attains the highest mean MRFS among the benchmarks considered. Experiments on 13 state-of-the-art MLLMs reveal a narrow accuracy range (31.4–42.1%), only modestly above the 20% chance baseline, exposing persistent limitations in multi-evidence reasoning. We trace these failures to two main bottlenecks: a retrieval deficit, where frame selectors fail to recover all necessary cues, and a fusion deficit, where models fail to combine them even when available, often collapsing onto a single frame. HERBench therefore establishes a principled benchmark for studying high-evidential-requirement video understanding, exposes substantial headroom beyond single-cue success, and points toward future improvements in retrieval-aware querying and distributed evidence fusion.
@@ 7.1 Track Ranking and Selection
To ensure robust performance in our benchmark’s highly dynamic environments—characterized by dense crowds and significant depth variations—we conducted a rigorous empirical evaluation of state-of-the-art detection and tracking pipelines. Specifically, we benchmarked three leading detectors (RF-DETR-L [38], YOLO-v11-x [17], and Grounding-DINO [24]) in combination with two widely adopted multi-object trackers (DeepSORT [46] and ByteTrack [58]). This evaluation was performed on a curated subset of the DanceTrack [42] benchmark, explicitly selected to mirror the crowd density, complex motion dynamics, and high inter-subject appearance diversity typical of our video corpus.
As detailed in Table 4, the combination of RF-DETR-L [38] and DeepSORT [46] yielded the highest tracking fidelity, achieving a HOTA of 47.1% and an IDF1 of 53.2%. Consequently, we adopted this configuration as the foundation of our pipeline. We utilize this detection-tracking stack by applying a high-recall RF-DETR [38] detector with a confidence threshold of 0.3 and a per-frame cap of 300 detections. Association within DeepSORT [46] uses a two-stage IoU matching: high-confidence detections (score >0.5>0.5) are matched with an IoU threshold of 0.7, followed by lower-confidence detections with a relaxed IoU threshold of 0.35.
HOTA / IDF1 (%) RF-DETR-L YOLO-v11-x Grounding-DINO
DeepSORT 47.1 / 53.2 43.3 / 51.1 42.9 / 50.6
ByteTracker 46.9 / 52.7 43.3 / 51.0 42.6 / 50.5
Table 4: Tracking pipeline benchmarking on a curated DanceTrack [42] subset mirroring HERBench’s dynamic characteristics. The combination of RF-DETR-L and DeepSORT achieved the highest scores and was selected for our data generation pipeline.
@@
Prior to ranking, we enforce a hard filter: we keep only the COCO [23] “person” class, require length $\geq 20$ frames, average area $\geq 5,500$ pixels, and require the track center to fall within the central safe region (frame cropped by 10% margins) in at least 5 frames.
Figure 6: Tracking, post-processing, and ranking pipeline. RF-DETR [38] detections are linked with DeepSORT [46] into raw person tracks, followed by outlier removal, gap interpolation, and Gaussian smoothing. A TrackRanker then scores and selects salient trajectories, which are passed to an MLLM descriptor module to generate temporally decoupled appearance (A) and behavior (B) cards that serve as the scaffold for downstream HERBench tasks.
@@
Track Selection as Noise Control.
Per video, we select the top 6–10 tracks according to the TrackRank composite score (Eq. 3). Post-hoc analysis of the selected tracks’ average detection confidence shows that they consistently fall within the top ${\sim}20\%$ most confident tracks per video, confirming that the ranking procedure implicitly filters out low-confidence, noise-prone trajectories before they can propagate errors into the question generation pipeline. The TrackRank selection process thus doubles as a noise control mechanism: by funneling only high-confidence, well-resolved trajectories into downstream task programming, it substantially mitigates the risk of identity switches, fragmented detections, or phantom tracks propagating into the ground-truth labels.
@@ 7.2 Decoupled Descriptor Generation
For each selected track, we generate disentangled descriptions using GPT-4o [33]. We sample 10-11 crops, reserving the first and last 20% of the trajectory for Appearance (A-cards) and the middle 60% for Behavior (B-cards). This ensures a temporal gap of at least 30 frames between appearance and behavior cues. An example of the resulting disentangled A- and B-cards for a single track is shown in Figure 7. We use the following prompt structure:
@@ 7.3 Spatial Operations and Region Definitions
For [RLPC], we also define rectangular ROIs (e.g., frame halves or specific zones). A track is counted as visiting an ROI if, at any frame, at least 50% of its bounding box area lies within the region (Intersection-Over-Box $\geq 0.5$). We count the unique track IDs that satisfy this predicate to derive people counts under spatial constraints. To absorb residual tracking noise (missed detections, fragmented tracks), multiple-choice options are reported as binned count ranges rather than exact integers. The bins are constructed so that the correct range spans approximately ±40% around the true count on average, ensuring that minor tracking errors do not invalidate the ground truth while still requiring models to perform meaningful spatial counting, see Figure 21 for example.
Duration computation ([MPDR]).
@@ 7.4 Scene Card Perturbations
We use TransNetV2 [41] for shot boundary detection. To calibrate its reliability on our video corpus, we manually reviewed shot boundaries on 30 videos used for [TSO] and [SVA] tasks (34% of videos contributing to these tasks). Comparing TransNetV2 [41] predictions against manual annotations yields F1 = 0.97, confirming that shot segmentation noise is negligible. For the [SVA] task, faithful scene cards are generated via an MLLM using the following prompt:
@@
To generate negative samples for [SVA], we prompt the model to modify faithful descriptions by altering 2-5 atomic details. The prompt constraints ensure:
@@
An example of a faithful scene card and its perturbed counterpart used for the [SVA] task is shown in Figure 8.
@@ 7.5 Corpus-Plausible Foil Generation
  • False Action Memory ([FAM]): We sample a “false” action by pairing an object present in the video with an action from the corpus that does not occur in the current video.
  • False Object Memory ([FOM]): We select an absent object from the corpus-wide index that is compatible with actions present in the video (e.g., if “cutting” occurs, “carrot” is a valid distractor if absent).
  • Action Counting ([AC]): Distractor counts are generated such that the correct count’s rank varies uniformly across options.
  • Action Sequence Integrity ([ASII]): We sample a 5-event ground-truth timeline. Distractors are generated using two perturbation functions: swap_mid (swapping two non-adjacent events) and rotate (shifting the sequence). Crucially, we verify against the event log that the perturbed timeline does not accidentally exist in the video.
@@ 7.6 Text-Only Bias Filtering Details
To suppress language priors, we apply a rigorous Text-Only Filtering stage. We discard any question correctly answered by $\geq 3$ of 4 blind LLMs (Qwen2-7B [36], Qwen2.5-7B [37], Llama-3-8B [30], and Vicuna-7B v1.5 [6]). This step rejects approximately 10% of candidates (e.g., questions answerable via object-color co-occurrence priors).
@@ 7.7 Human Verification Protocol
Experts conduct verification on a stratified 15% sample of instantiated questions to audit whether the construction pipeline preserves the intended evidential constraints. The audit focuses on three properties: (i) minimum frame-set compliance, i.e. confirming that the item requires at least three distinct frames; (ii) uniqueness of the answer, i.e. confirming the existence of a single objective ground-truth answer; and (iii) descriptor disentanglement, i.e. verifying that A-cards and B-cards do not leak information from one another. Items that violate any of these conditions are rejected. This process resulted in a 17.8% rejection rate.
7.8 Human Validation and Oracle-Frame Study
Task Full video (%) Oracle frames (%)
[TSO] 91.7 93.8
[MPDR] 88.3 97.0
[ASII] 86.7 95.8
[AGBI] 95.8 98.1
[AGAR] 95.0 98.3
[AGLT] 92.5 97.4
[FAM] 84.2 92.8
[SVA] 84.2 94.8
[FOM] 87.5 96.3
[MEGL] 85.8 95.4
[AC] 84.2 97.7
[RLPC] 90.0 90.9
Overall 88.8 95.7
Table 5: Per-task human accuracy. Accuracy of human annotators in the full-video and oracle-frame settings across all HERBench tasks. In the full-video setting, annotators answer with unrestricted video access and free scrubbing. In the oracle-frame setting, annotators answer using only the curated oracle frame-set, without access to the source video.
To complement the construction-time audit above, we conducted two human studies that assess HERBench from two complementary perspectives: overall question answerability and oracle-evidence sufficiency. Each study used a separate group of 6 annotators.
Study design.
In the full-video setting, annotators answered a shared set of 240 questions spanning all 12 tasks (20 questions per task) with unrestricted video access and free scrubbing. In the oracle-frame setting, annotators answered 2,160 questions using only the curated oracle frame-set provided by the benchmark construction pipeline, without access to the source video. Each oracle item was presented in the same format used by the oracle-based analysis in the main paper, namely the curated evidence frames together with distractor frames.
Results.
Table 5 reports the per-task accuracies. In the full-video setting, annotators achieved 88.8% accuracy overall, with substantial inter-annotator agreement (Fleiss’ $\kappa=0.74$), indicating that the benchmark remains highly answerable for humans despite its high evidential demand. In the oracle-frame setting, annotators achieved 95.7% accuracy overall, showing that the curated oracle frame-set is generally sufficient to resolve the question without access to the full temporal context. The largest improvements appear in tasks such as [AC], [SVA], and [MEGL], where the answer depends on sparse or temporally localized evidence.
Benchmark cleanup.
After each oracle-frame response, annotators were shown the ground-truth answer and invited to flag problematic items. This process surfaced three types of issues: mis-indexed evidence frames, incorrect ground-truth labels, and genuinely ambiguous items. Specifically, annotators flagged 42/2160 items (1.9%) for evidence mis-indexing, 18/2160 (0.8%) for incorrect ground truth, and 58/2160 (2.7%) as ambiguous. All flagged items were subsequently corrected or removed from the final release.